What is the perimeter of a rectangle or garden that is 16 units wide and has the same area as a rectangular garden that is 10 units wide and 32 units long?
Answer by rothauserc(4718)

The area of the second rectangular garden is
:
Area = 10 * 32 = 320 square units
:
Let l be the length of the first rectangular garden
:
l * 16 = 320
:
l = 20
:
*********************************************************
Perimeter of first rectangle = 2 * 20 + 2 * 16 = 72 units
*********************************************************
: